# LLMs Can’t Plan, But Can Help Planning in LLM-Modulo Frameworks

### 06 Feb 2024 v2

@@ 2.3 Analyzing Claims to the Contrary in the Literature
The guarantees–if any–are coming in terms of soundness of the external verifier. The one clear reasoning problem used in the ToT paper is the 24 puzzle–for which the external verifier can be easily implemented in terms of arithmetic operations (thankfully not done by the numerically challenged LLM!). Here, our experiments show that LLM’s own criticisms are often quite off the mark. [^8] Note that we can do this check easily because of the formal specification of correctness. For the “improving writing task” also used in ToT, there are no formal quality metrics and so it is hard to say anything concrete about the critiques of the LLM. Because the 24 puzzle’s solutions can be verified by simple arithmetic operations, readers don’t quite realize that the framework relies on an external verifier. In general though, the verifier may be more complex and can involve substantial work (you can substitute a simulator for the verifier–but someone has to write that simulator too!)
In general planning problems, one way to provide an external verifier is to (a) write a domain model (e.g. in PDDL) and (b) feed it to an off-the-shelf model-based verifier like VAL (c.f. Howey et al. (2004)).
@@ 3 LLM-Modulo Framework for Robust Planning
While Section 2 questions the claims that LLMs are capable of planning/reasoning by themselves, it is certainly not meant to imply that LLMs don’t have any constructive roles to play in solving planning/reasoning tasks. On the contrary, as discussed in the Introduction, their uncanny ability to generate ideas/potential candidate solutions–albeit with no guarantees about those guesses–can be valuable in the generate-test-critique setups in conjunction with either model-based verifiers or expert humans in the loop. Accordingly, we propose a general “LLM-Modulo” framework [^9] The name LLM-Modulo is inspired by the SAT-Modulo theories Nieuwenhuis & Oliveras (2006). While we believe that versions of such an architecture can be of use in a wide variety of planning or reasoning tasks, for the sake of concreteness, we will focus on planning tasks, especially of the type studied in the automated planning community Ghallab et al. (2004).
@@
Figure 3 gives a schematic of the LLM-Modulo Framework, as we envision it. As can be seen readily, the underlying architecture is a Generate-Test-Critique loop, with the LLM generating candidate plans and a bank of critics critiquing the candidate. The loop starts with the LLM getting the problem specification and generating its first plan candidate. [^10] Although we focus on planning from scratch, it is easy to accommodate replanning scenarios, where the loop starts with an externally supplied candidate plan. Note that the plans an LLM helps generate in this architecture have soundness guarantees because of the external sound critics. This means that plans coming out of such an architecture will constitute a better corpus of synthetic data for any fine tuning phase carried out to improve/customize the LLM’s generation capability.
@@
Design Considerations: Before going into the details about the framework and its various modules, it is worth noting some design decisions underlying the proposed architecture. We start by noting that the LLM-Modulo architecture is a “Generate-Test” one that involves LLMs interacting with the external critics rather than solvers. This is a deliberate decision–as this allows the LLM to guess/generate candidates to satisfy the critics, as against dealing with the expressiveness and search complexity issues of the solvers. Secondly, the framework explicitly recognizes that the LLMs can generate approximate ideas not just about plan candidates, but domain models, problem reduction strategies, and refinements to the problem specification. The framework also recognizes that LLMs are good at format/syntax changes. Accordingly, the framework leverages all these abilities of LLMs, letting them play multiple roles in planning. Finally, the architecture carefully circumscribes the human’s role–domain experts interact with the LLM to tease out the models used by (some of) the critics, while end users take part in refinining any incomplete problem specification in concert with the LLM. A notable, and deliberate, absence is human’s involvement in the inner loop of planning–e.g. with iterative prompting. In addition to posing an infeasible burden on the human’s time for complex planning problems, such iterative prompting strategies are notorious for their Clever Hans effect cle .
3.1 Critics
@@
In the LLM-Modulo framework, critics can evaluate LLM-generated candidates for a planning/reasoning problem over both hard and soft constraints. Hard constraints refer to correctness verification which can include causal correctness, timeline correctness, resource constraint correctness, etc. For PDDL planning problems, the hard critic can be based on VAL Howey et al. (2004), that works off of a model (which itself can be acquired with the help of the LLM Guan et al. (2023). On the other hand, soft constraints can include more abstract notions of correctness such as style, explicability, preference conformance, etc. As discussed in Section 2.3, while LLMs cannot take on the role of hard critics, they can help simulate some aspects of the role of soft critics. So our framework does allow for style critics be possibly based on LLMs (e.g Verma et al. (2024)). We reiterate that the soundness of the LLM-modulo framework is inherited from the soundness of the critics.
@@
The bank of critics–hard (model-based) as well as soft (possibly LLM-based) evaluate the current plan candidate to evaluate its fitness/acceptability. If at least all the hard critics sign off on the current candidate, then that is considered a valid solution to be returned to the end-user or the executor. When a critic finds the current plan candidate to be unsatisfactory, it can provide varying levels of feedback, ranging from “No, try again” to “No, try again, here is one thing wrong with the current plan” to “No, try again, here are all the things wrong with the current plan. These critiques are all pooled at the Backprompt Controller (see Section 3.2)
3.1.1 LLMs as Reformulators
@@
One interesting challenge is that many of the symbolic model-based verifiers tend to be operating on specialized formal representations. Given a central candidate plan (e.g. a mission plan), these critics need translations of that candidate into their representations. This is the role of the reformulator module attached to individual critics. These reformulator modules can be supported to large extent by LLMs, given that one thing LLMs are very good at is format change across different syntactic representations, Olmo et al. (2021). Indeed, as discussed in the Introduction, some approaches to combine LLMs with external symbolic solvers just use LLMs as reformulators for these solvers Liu et al. (2023); Pan et al. (2023). Our discussion of LLM-Modulo framework should make it clear that syntax reformulation alone is a severely limited role for LLMs!
@@ 3.2 Backprompt (Meta) Controller
The critiques from the various critics are pooled together by the Meta (Backprompt) Controller, which passes a processed version of them to the LLM as the next iterative prompt to elicit the next guess. This is especially required in the presence of a mix of soft and hard critics, where the Meta Controller can assume the responsibility of compiling the critiques into a consistent feedback to process.
@@
The processing steps taken in the controller can range from simple round-robin selection of prompts to generating a summarized prompt (with LLM help)111111 Such summarization is a reasonable strategy as the back prompts will not be treated as hard constraints by LLMs anyway. to employ a prompt diversification strategy to elicit the next candidate from a different part of the implicit search space (akin effectively to the strategy used in systems such as Tree of Thoughts prompting Yao et al. (2023a), as discussed in 2.3).
3.3 Fine Tuning & Synthetic Data
Once the LLM-Modulo framework “solves” a planning instance, the solution can then be added to a synthetic data corpus (step 6 in Figure 3), which is intermittently used to fine tune the LLM (step 7), so its future plan candidate guesses improve.
Such fine tuning on task-specific data has been a popular way to get LLMs to improve their performance on reasoning/planning tasks. For example, fine tune the LLM on blocks world planning problem-solution pairs to improve their performance in guessing solutions for blocks world instances Pallagani et al. (2023). While fine tuning still doesn’t guarantee correctness of the generated solutions, it might improve the chances that LLM guesses candidates that are closer to being vetted by the bank of critics.
One important question is where this additional data for fine tuning comes from. A tempting idea is to have the LLM itself generate this additional data, and improve it by self-critiquing/verification, before fine-tuning itself on the data. This EM-like approach unfortunately will not work given that LLMs can’t verify their own solutions (see Section 2.2). In the past, this meant that the only reliable way to generate synthetic data is to use external plan generators–for example, use a classical planner like FF Hoffmann & Nebel (2001) to solve a host of blocks world instances and then use those solutions to fine-tune the LLM. The LLM-Modulo framework, in contrast, provides an alternative way of using an LLM-based framework to generate synthetic data that is guaranteed correct.
3.4 Specification Refinement & Model Acquisition (Semi-automated)
@@
As mentioned earlier, we avoid having humans involved in iteratively prompting LLMs–as this can be an infeasibly time-consuming activity for them. Instead, we let automated verifiers, either model-based or LLM-supported, to manage the plan critiquing process. The framework does depend on humans for “once per domain” and “once per problem” interactions. In the former category, human domain experts can play a role in acquiring the domain model with the help of the LLM. Examples of such interaction include teasing out PDDL planning models from the LLMs with the help of human expert curation (top left in Figure 3). The idea here is that the traditional domain model acquisition task (e.g. sim (2001)) is significantly made easier by having the LLMs help with ideas regarding various pieces of the domain model (e.g., actions, their preconditions and effects) and letting humans sign off/critique the resulting model. Once the model is acquired this way, it can be used by correctness verifiers such as VAL Howey et al. (2004); Guan et al. (2023). Often the planning problems in real world situations are specified incompletely, leaving it to the human commonsense to refine the specification. This brings up a second role for humans–this time end users (bottom left in Figure 3–in collaboratively refining the specification with the help of LLMs (similar to the way done in Xie et al. (2023); Liu et al. (2023)).
3.5 Summary of Roles of LLMs in the LLM-Modulo Framework
@@
It is worth summarizing the multiple roles the LLM plays in the LLM-Modulo architecture. The most prominent, of course, is its role in “guessing” the candidate plans (step 2 in Figure 3) in response to problem specification and iterative back prompting from the bank of critics (Step 5). Second the LLM plays a role in converting the guessed plan candidate into specialized representations used by the various critics (e.g., the time-line view, the causal link view etc.). This role leverages the fact that LLMs are very good at format conversion (c.f. Olmo et al. (2021)) Third, the LLM plays a role in helping the end user flesh out the incomplete problem specification to begin with (Step 1 in Figure 3). Finally, the LLM plays a role in helping the domain expert tease out and refine the domain models used by the various model-based critics Guan et al. (2023); Kwon et al. (2022).

### 12 Jun 2024 v3

@@ 2.3 Analyzing Claims to the Contrary in the Literature
The guarantees–if any–are coming in terms of soundness of the external verifier. The one clear reasoning problem used in the ToT paper is the 24 puzzle–for which the external verifier can be easily implemented in terms of arithmetic operations (thankfully not done by the numerically challenged LLM!). Here, our experiments show that the LLM’s own criticisms are often quite off the mark. [^6] Note that we can do this check easily because of the formal specification of correctness. For the “improving writing task” also used in ToT, there are no formal quality metrics and so it is hard to say anything concrete about the critiques of the LLM. Because the 24 puzzle’s solutions can be verified by simple arithmetic operations, it is trivial to implement an external verifier for these problems. In general though, the verifier may be more complex and can involve substantial work (you can substitute a simulator for the verifier–but someone has to write that simulator too!).
LLMs as Approximate Knowledge Sources: The fact that LLMs are often good at extracting planning knowledge can indeed be gainfully leveraged. As shown in recent works (Guan et al., 2023), LLMs can be a rich source of approximate models of world/domain dynamics and user preferences, as long as the humans (and any specialized critics) in the loop verify and refine those models, and give them over to model-based solvers. This way of using LLMs has the advantage that the humans need only be present when the dynamics/preference model is being teased out and refined, and the actual planning after that can be left to sounder planning frameworks with correctness guarantees, such as the LLM-Modulo framework we propose.
Such an overall approach has striking similarities to knowledge-based AI systems of yore, with LLMs effectively replacing the “knowledge engineer” (see Figure 2). Given the rather quixotic and dogmatic shift of AI away from approaches that accept domain knowledge from human experts that can be termed “Polanyi’s Revenge” (c.f. (Kambhampati, 2021)), this new trend of using LLMs as knowledge sources can be viewed as a form of avenging Polanyi’s revenge! Indeed, LLMs make it easy to get problem-specific knowledge as long as we are willing to relax the correctness requirements of that knowledge. In contrast to the old knowledge engineering approaches, LLMs offer this without making it look like we are inconveniencing any specific human (we are, instead, just leveraging everything humans told each other on the Web!). So the million dollar question for reasoning tasks is: “how would you do robust planning if you have some doddering know-it-all ready to give you any kind of knowledge?” The LLM-Modulo Framework is a principled method for tackling this challenge.
@@ 3 LLM-Modulo Framework for Robust Planning
While Section 2 questions the claims that LLMs are capable of planning/reasoning by themselves, it is certainly not meant to imply that LLMs don’t have any constructive roles to play in solving planning/reasoning tasks. On the contrary, as discussed in the Introduction, their uncanny ability to generate ideas/potential candidate solutions–albeit with no guarantees about those guesses–can be valuable in the generate-test-critique setups in conjunction with either model-based verifiers or expert humans in the loop. Accordingly, we propose a general “LLM-Modulo” framework [^7] The name LLM-Modulo is inspired by the SAT-Modulo theories (Nieuwenhuis & Oliveras, 2006).. While we believe that versions of such an architecture can be of use in a wide variety of planning or reasoning tasks, for the sake of concreteness, we will focus on planning tasks, especially of the type studied in the automated planning community (Ghallab et al., 2004).
@@
Figure 3 gives a schematic of the LLM-Modulo Framework, as we envision it. As can be seen readily, the underlying architecture is a Generate-Test-Critique loop, with the LLM generating candidate plans and a bank of critics critiquing the candidate. The loop starts with the LLM getting the problem specification and generating its first plan candidate. [^8] Although we focus on planning from scratch, it is easy to accommodate replanning scenarios, where the loop starts with an externally supplied candidate plan. Note that the plans an LLM helps generate in this architecture have soundness guarantees because of the external sound critics. This means that plans coming out of such an compound system will constitute a better corpus of synthetic data for any fine tuning phase carried out to improve/customize the LLM’s generation capability. The completeness of the system depends on the LLM’s ability to generate all potentially relevant candidates.
Design Choices: Before going into the details about the framework and its various modules, it is worth noting some design decisions underlying the proposed architecture. We start by noting that the LLM-Modulo architecture is a “Generate-Test” one that involves LLMs interacting with the external critics/verifiers rather than a LLMs being just front-ends to external solvers. This is a deliberate decision–as this allows the LLM to guess/generate candidates to satisfy the critics, as against dealing with the expressiveness and search complexity issues of the solvers. The critics/verifiers also are also more naturally composable than solvers/planners. As we shall see, we do allow for constructive critics which can be based on solvers, and provide suggestions on specific ways of extending/modifying the candidate plans.
@@
Secondly, the framework explicitly recognizes that the LLMs can generate approximate ideas not just about plan candidates, but domain models, problem reduction strategies, and refinements to the problem specification. The framework also recognizes that LLMs are good at format/syntax changes. Accordingly, the framework leverages all these abilities of LLMs, letting them play multiple roles in planning. Finally, the architecture carefully circumscribes the human’s role–domain experts interact with the LLM to tease out the models used by (some of) the critics, while end users take part in refining any incomplete problem specification in concert with the LLM. A notable, and deliberate, absence is human’s involvement in the inner loop of planning–e.g. with iterative prompting. In addition to posing an infeasible burden on the human’s time for complex planning problems, such iterative prompting strategies are notorious for their Clever Hans effect (cle, ).
3.1 Critics/Verifers
@@
In the LLM-Modulo framework, critics can evaluate LLM-generated candidates for a planning/reasoning problem over both hard and soft (style) constraints. Hard constraints refer to correctness verification which can include causal correctness, timeline correctness, resource constraint correctness as well as unit tests. For PDDL planning problems, the hard critic can be based on VAL (Howey et al., 2004), that works off of a model (which itself can be acquired with the help of the LLM (Guan et al., 2023). It is worth noting that the critics don’t always have to be declarative model-based ones, and can be simulators. Just as LLMs can help humans in coming up with models, they can also help in writing procedural simulators, as seems to be done in systems like Voyager (Wang et al., 2023a).
On the other hand, soft constraints can include more abstract notions of good form such as style, explicability, preference conformance, etc. As discussed in Section 2.3, while LLMs cannot take on the role of hard critics with soundness guarantees, [^9] If we don’t insist on soundness guarantees, then it is, in principle, possible to train LLMs discriminatively to learn to verify plans; see (Arora & Kambhampati, 2023). they can help simulate some aspects of the role of soft (style) critics. So our framework does allow for style critics be possibly based on LLMs. For example, in (Verma et al., 2024b) we discuss how LLMs can act as a human proxy to evaluate plans in terms of how they would be perceived by humans in the loop. Additionally, in (Guan et al., 2024), we show how Vision-Language Models (VLMs) can be leveraged to critique the style of robot behaviors in terms of their adherence to the soft common-sense preferences of the humans in the loop. We reiterate that the soundness of the LLM-modulo framework is inherited from the soundness of the correctness (hard) critics.
@@
The bank of critics–hard (model-based) as well as soft (possibly LLM-based) evaluate the current plan candidate to evaluate its fitness/acceptability. If at least all the hard critics sign off on the current candidate, then that is considered a valid solution to be returned to the end-user or the executor. When a critic finds the current plan candidate to be unsatisfactory, it can provide varying levels of feedback, ranging from “No, try again” to “No, try again, here is one thing wrong with the current plan” to “No, try again, here are all the things wrong with the current plan.” More importantly, the critics can be constructive, and offer alternatives plan/subplan suggestions. One way of obtaining such constructive critics is to base them on partial planners–operating either on the models themselves or their relaxations (Bryce & Kambhampati, 2007). These critiques are all pooled at the Meta (Backprompt) Controller (see Section 3.2)
@@
LLMs as Reformulators: One interesting challenge is that many of the symbolic model-based verifiers tend to be operating on specialized formal representations. Given a central candidate plan (e.g. a mission plan), these critics need translations of that candidate into their representations. This is the role of the reformulator module attached to individual critics. These reformulator modules can be supported to a large extent by LLMs, given that one thing LLMs are very good at is format change across different syntactic representations (Olmo et al., 2021). Indeed, as discussed in Section 1, some approaches to combine LLMs with external symbolic solvers just use LLMs as reformulators for these solvers (Liu et al., 2023; Pan et al., 2023). It is worth noting that the syntax conversion itself can be helped with a nested LLM-Modulo framework–where the syntactic correctness of the conversion is checked by syntax critics. We will have occasion to illustrate this in the context of our preliminary work on LLM-Modulo frameworks for travel planning discussed in Section 4. Our discussion of LLM-Modulo framework should make it clear that syntax reformulation alone is a severely limited role for LLMs!
@@ 3.2 Backprompt (Meta) Controller
The critiques from the various critics are pooled together by the Meta (Backprompt) Controller, which passes a processed version of them to the LLM as the next iterative prompt to elicit the next guess. This is especially required in the presence of a mix of soft and hard critics, where the Meta Controller can assume the responsibility of compiling the critiques into a consistent feedback to send to the LLM.
@@
The processing in the controller can range from (i) simple round-robin selection of prompts to (ii) generating a summarized prompt (with LLM help) to (iii) employing a prompt diversification strategy to elicit the next candidate from a different part of the implicit search space. This last strategy helps increase the completeness of the LLM candidate generation, and may involve domain/task-specific knowledge (see the discussion of Tree of Thoughts in Section 2.3).
3.3 Specification Refinement & Critic/Model Acquisition
@@
As mentioned earlier, we avoid having humans involved in iteratively prompting LLMs–as this can be an infeasibly time-consuming activity for them. Instead, we let automated verifiers, either model-based or LLM-supported, to manage the plan critiquing process. The framework does depend on humans for “once per domain” and “once per problem” interactions. In the former category, human domain experts can play a role in acquiring the domain model with the help of the LLM. Examples of such interaction include teasing out PDDL planning models from the LLMs with the help of human expert curation (top left in Figure 3). An example of this is our work in (Guan et al., 2023). The idea here is that the traditional domain model acquisition task (e.g. (Simpson et al., 2001)) is significantly made easier by having the LLMs help with ideas regarding various pieces of the domain model (e.g., actions, their preconditions and effects) and letting humans sign off/critique the resulting model. Once the model is acquired this way, it can be used by correctness verifiers such as VAL (Howey et al., 2004; Guan et al., 2023). Often the planning problems in real world situations are specified incompletely, leaving it to the human commonsense to refine the specification. This brings up a second role for humans–this time end users (bottom left in Figure 3–in collaboratively refining the specification with the help of LLMs (similar to the way done in (Xie et al., 2023; Liu et al., 2023)).
3.4 Summary of LLM Roles in LLM-Modulo
@@
It is worth summarizing the multiple roles the LLM plays in the LLM-Modulo architecture. The most prominent, of course, is its role in “guessing” the candidate plans (step 2 in Figure 3) in response to problem specification and iterative back prompting from the bank of critics (Step 5). Second, the LLM plays a role in converting the guessed plan candidate into specialized representations used by the various critics (e.g., the time-line view, the causal link view etc.). This role leverages the fact that LLMs are very good at format conversion (c.f. (Olmo et al., 2021)). Third, the LLM plays a role in helping the end user flesh out the incomplete problem specification to begin with (Step 1 in Figure 3). Finally, the LLM plays a role in helping the domain expert tease out and refine the domain models used by the various model-based critics (Guan et al., 2023; Kwon et al., 2022), or help “implement” procedural critics (such as those checking syntactic constraints). As a broad approximate source of knowledge, the LLM can also help enumerate the list of potential critics needed to validate the candidate plans (once again with a human in the loop).
Figure 4: LLM Modulo Framework adapted for Travel Planning
Figure 5: Final Pass rates of models across LLM Modulo Iterations
3.5 Can LLM-Modulo Frameworks Pay Their Way?
Let’s address the elephant in the room: Given that formal model-based planning systems already exist (Ghallab et al., 2004), is LLM-Modulo framework for planning more than a gratuitous attempt to shoe-horn (the currently popular) LLMs to solve planning problems? Indeed, when the underlying problem is actually solvable by such combinatorial solvers, it can be orders of magnitue more resource efficient to use them. [^10] Not surprisingly, automated programming, the one community that certainly doesn’t have the luxury of a ready-made “solver,” have stuck to LLM-Modulo style approaches. Compared to a planner that is guaranteed to be correct in a narrow set of domains, LLMs may likely be good at generating plausible (but not guaranteed to be correct) plan heuristics/suggestions in many more scenarios. Thus, unlike the traditional planning architectures studied in AI (Ghallab et al., 2004), which put a priori constraints on the expressiveness of the problems that can be posed to the planner (to wit, the different expressiveness levels of the PDDL specification (McDermott et al., 1998)), the LLM-Modulo architecture puts no such restrictions. In this sense, it is more representative of real-world planning problems such as those in NASA mission planning, where the different critics–human and automated–are at best able to give “no objection” certificates for the candidate plans under consideration, clearing it from their perspective. (Indeed, both deep space network planning and mars rover task planning are done via a collective human blackboard. (Johnston et al., 2014; Bresina et al., 2004).) Note that this is starkly different from just sending an unvetted plan out to execution (as would be the case if we have LLMs operate in autonomous mode to guess plans). Generalizing planning and reasoning frameworks this way is consistent with the Doyle & Patil’s call to the Knowledge Representation community of yore (Doyle & Patil, 1991), as well as our own call for model-lite planning (Kambhampati, 2007).
4 Two Case Studies of LLM-Modulo
We have applied the LLM-Modulo framework to classical planning domains (as reported in (Valmeekam et al., 2023c)) and to a recent travel planning benchmark (as reported in (Gundawar et al., 2024)). In the former case, the results (presented in Section 5.2 and Table 4 of (Valmeekam et al., 2023c)) show that with back prompting from VAL (Howey et al., 2004) acting as the external verifier and critic, LLM performance in Blocks World improves to 82% within 15 back prompting rounds, while in Logistics, it improves to 70%. LLM-Modulo doesn’t help as much in an obfuscated version of blocks world called Mystery BW, reaching about 10% accuracy. This should be expected because the LLMs have difficulty generating plausible candidate plans for this domain (note that even here, if a plan is returned, it must have passed muster with VAL, and is thus guaranteed correct by its model).
For the travel planning case study, we used a benchmark proposed in (Xie et al., 2024), which involves a rich mix of travel constraints presented in flexible natural language form. Our preliminary results on adapting LLM-Modulo framework to this benchmark are reported in (Gundawar et al., 2024). The benchmark’s authors test LLMs across a variety of prompt engineering techniques including Chain of Thought and ReAct, reporting that–on GPT-3.5-Turbo–the current best strategies only manage a startlingly low 0.7% performance rate! We adapted the LLM-Modulo framework to this benchmark by operationalizing their hard constraints (such as the budget constraint set by the user) or common-sense constraints (such as suggesting diverse attractions to visit) as critics as shown in Figure 4. Our preliminary results show (see Figure 5; additional results in (Gundawar et al., 2024)) that LLM-Modulo based agentification with automated critics in the loop significantly improves the performance (6x of baselines) even with a limit of 10 back prompting cycles, and weaker models such as GPT-3.5-turbo. Furthermore, we also find that LLMs can successfully implement functions corresponding to hard critics and several common-sense critics. Finally, LLMs reliably play the role of reformatter as well, converting free form travel plans into structured plans parseable by the critics for backprompts or plan evaluation. One interesting observation about this domain is that we were able to use the LLM itself to enumerate the type of critics needed to validate the plan (with light human supervision).